OpenStudy (anonymous):

Need help to solve this differential equation \frac{dy}{dx} + \frac{2}{\tan(2x)} \times y = \cos(2x)

OpenStudy (dan815):

\[\frac{dy}{dx} + \frac{2}{\tan(2x)} \times y = \cos(2x)\]

OpenStudy (dan815):